he spot rates are

Time to maturity Spot rate
1 year 3.00%
2 years 4.00%
3 years 4.70%
4 years 5.20%

What is the yield to maturity of a 4-year bond with a six percent annual coupon? Answer to the nearest 0.01%, leaving off the % sign.
Answer #1

Face Value=100
Coupons=Coupon rate*Face (or Par) value=6%*100=6
Price of the bond=6/1.03+6/1.04^2+6/1.047^3+6/1.052^4+100/1.052^4=103.1454707

Using financial calculator
N=4
PMT=6%*100=6
FV=100
PV=-103.1454707
CPT I/Y=5.111%

Hence, yield to maturity=5.111%